OpenStudy (anonymous):

Lyle has a thirteen-cup box of cereal. If one serving is three-fourths cup, how many servings are in the box? 9 and three-fourths thirteen and three-fourths fifteen and two-thirds seventeen and one-third

OpenStudy (anonymous):

help

jimthompson5910 (jim_thompson5910):

1 serving = 3/4 of a cup

jimthompson5910 (jim_thompson5910):

we need to get the right side of 3/4 of a cup to 13 cups, how do we do this?

OpenStudy (anonymous):

idk

jimthompson5910 (jim_thompson5910):

(3/4)x = 13 3x = 13*4 3x = 52 x = 52/3 we need to multiply both sides by 52/3

jimthompson5910 (jim_thompson5910):

1 serving = 3/4 of a cup (52/3)*1 serving = (52/3)*3/4 of a cup 52/3 servings = 13 cups (51+1)/3 servings = 13 cups 51/3+1/3 servings = 13 cups 17+1/3 servings = 13 cups 17 & 1/3 servings = 13 cups

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it would be 13 3/4

jimthompson5910 (jim_thompson5910):

no i posted the answer in the last line

OpenStudy (anonymous):

17 and 1/3

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

np